# A Connection Between Covariate Adjustment and Stratified Randomization in Randomized Clinical Trials

### v2

@@ 6 Discussion
With focus on statistical efficiency, this article highlights a connection between covariate adjustment and stratified randomization with important practical implications. Specifically, stratified randomization can be regarded as a special form of covariate adjustment that makes optimal use of stratum information. Covariate adjustment is more general and able to incorporate information from all baseline covariates associated with clinical outcomes, although its effectiveness depends on the appropriateness of the method. In the analysis of trial data following stratified randomization, adjusting for the stratification factor only provides no efficiency benefit, and the key to efficient estimation is incorporating new prognostic information from other covariates. In designing a trial with stratified randomization, it is not essential to include all important covariates in the stratification, because their prognostic information can be incorporated through covariate adjustment. These findings are supported by both theoretical and empirical results.
@@
Appendix: Proofs
Proof of Equation (3)
@@
$\sigma_{\text{st}}^{2}(b)-\sigma^{2}(b)=-\pi(1-\pi)\|\Pi(b_{\text{opt}}-b|\mathcal{C})\|_{2}^{2}=-\pi(1-\pi)\|\Pi(b-b_{\text{opt}}|\mathcal{C})\|_{2}^{2}.$ (A.1)
@@
$\sigma^{2}(b)-\sigma^{2}(b_{\text{opt}})=\pi(1-\pi)\|b-b_{\text{opt}}\|_{2}^{2}.$ (A.2)
Adding (A.1) and (A.2) yields

### v3

@@ 6 Discussion
With focus on statistical efficiency, this article highlights a connection between covariate adjustment and stratified randomization with important statistical and practical implications. Specifically, stratified randomization can be regarded as a special form of covariate adjustment that makes optimal use of stratum information. Covariate adjustment is more general and able to incorporate information from all baseline covariates associated with clinical outcomes, although its effectiveness depends on the appropriateness of the method. In designing a trial with stratified randomization, it is not essential to include all important covariates in the stratification, because their prognostic information can be incorporated through covariate adjustment. In the analysis of trial data following stratified randomization, adjusting for the stratification factor only provides no efficiency benefit, and the key to efficient estimation is incorporating information from all important covariates. These findings are supported by both theoretical and empirical results.
@@
While we have adopted a superpopulation perspective in this article, it is worthwhile to relate our findings to recent developments in covariate adjustment and CAR in the literature on randomization-based causal inference 37, 38, 39, where study subjects (together with their potential outcomes and covariate values) are considered fixed and randomized treatment assignment is the only source of randomness. These authors study unadjusted, stratified and ANCOVA-type estimators of the sample average treatment effect but not general augmented estimators. The CAR mechanisms they consider are not as relevant to clinical trials as the stratified randomization mechanism considered in this article. Specifically, the rerandomization design considered by Li and Ding 38 is infeasible for clinical trials as it assumes that covariate data are available for all patients before randomization occurs, and the stratified randomization design in Miratrix et al. 37 and Liu and Yang 39 assumes that each stratum is randomized as one block, which is rarely the case in clinical trials. Nonetheless, there are interesting similarities between their findings and ours. For example, Theorem 4 of Miratrix et al. 37, a finite-sample result indicating that the variance of the stratified estimator is similar but slightly lower under stratified versus simple randomization, is consistent with our asymptotic theory and helps explain the empirical observation that stratified randomization seems slightly more effective than stratified analysis at $n=200$. The observation in Section 4.1.1 of Li and Ding 38, that rerandomization does not change the optimally adjusted estimator or its asymptotic distribution, is consistent with our finding in Section 3.2.1 that $\widehat{\delta}_{\text{aug}}(b_{\text{opt}})$ is optimal under both simple and stratified randomization with the same asymptotic variance. The asymptotic near equivalence between rerandomization and optimal adjustment, noted in Section 4.1.2 of Li and Ding 38, is similar in spirit to our asymptotic equivalence result in Section 3.2.3.
Appendix A: Regression Estimators
A regression (or G-computation) estimator of $\delta$ may be obtained by fitting a standard regression model:
$\operatorname{E}(Y|A,\boldsymbol{V})=h\left((1,A,\boldsymbol{V}^{\text{T}},A\boldsymbol{V}^{\text{T}})\mbox{\boldmath${\beta}$}\right),$ (A.1)
where $\boldsymbol{V}$ is a vector-valued function of $\boldsymbol{W}$, $h$ a specified inverse link function, and ${\beta}$ an unknown parameter vector. The interaction term in model (A.1) is optional in practice but necessary in the present discussion. Suppose ${\beta}$ is estimated by solving the estimating equation
$\sum_{i=1}^{n}\left\{Y_{i}-h(1,A_{i},\boldsymbol{V}_{i}^{\text{T}},A_{i}\boldsymbol{V}_{i}^{\text{T}})\mbox{\boldmath${\beta}$}\right\}(1,A_{i},\boldsymbol{V}_{i}^{\text{T}},A_{i}\boldsymbol{V}_{i}^{\text{T}})^{\text{T}}=\boldsymbol{0}.$ (A.2)
[…]
Examples of the regression estimator include the analysis of covariance (ANCOVA) estimator for continuous outcomes 3 and the standardized logistic regression estimator for binary outcomes 6. It is worth noting that, whether model (A.1) is correct or not, $\widehat{\delta}_{\text{reg}}$ is consistent for $\delta$ and asymptotically normal with asymptotic variance $\operatorname{Var}\{\psi(\boldsymbol{O})-(A-\pi)b_{\text{reg}}(\boldsymbol{W})\}$, where
[…]
Appendix B: Proofs
Proof of Proposition 1
@@
$\sigma_{\text{st}}^{2}(b)-\sigma^{2}(b)=-\pi(1-\pi)\|\Pi(b_{\text{opt}}-b|\mathcal{C})\|_{2}^{2}=-\pi(1-\pi)\|\Pi(b-b_{\text{opt}}|\mathcal{C})\|_{2}^{2}.$ (B.1)
@@
$\sigma^{2}(b)-\sigma^{2}(b_{\text{opt}})=\pi(1-\pi)\|b-b_{\text{opt}}\|_{2}^{2}.$ (B.2)
Adding (B.1) and (B.2) yields